10 Books Fit In A 1ft Container And There Are Currently 10,000 books what is the volume the must fit all the cubic containers
1
Expert's answer
2012-02-09T08:33:29-0500
If we have 10,000 books then we need 10000 / 10 = 1000 1ft containers,
but 1000 = 10 * 10 * 10 that means that we should have 10 ft length bug cubic container and his volume is 10*10*10 = 1000